# First Responders’ Perceptions of Semantic Information for Situational Awareness in Robot-Assisted Emergency Response

### v1

@@ Abstract
To the best of our knowledge, this study offers novel insights by being one of the first to directly survey FRs on semantic-based SA in a cross-national context. It reveals the types of semantic information most valued in the field—such as object identity, spatial relationships, and risk context—and connects these preferences to the respondents’ roles, experience, and education levels. The findings also expose a critical gap between lab-based robotics capabilities and the realities of field deployment, highlighting the need for more meaningful collaboration between FRs and robotics researchers. These insights contribute to the development of more user-aligned and situationally aware robotic systems for emergency response.
@@ I Introduction
At the same time, robotic technologies — such as drones, ground vehicles, and other autonomous systems — are becoming increasingly accessible and capable, due to rapid advances in perception, planning, and control. Commercial applications are expanding, yet these developments have not fully translated into effective tools for FRs operating in real-world extreme conditions. Despite growing interest, a significant gap persists between laboratory robotics and field deployments. Challenges such as rugged terrain, environmental hazards (e.g., heat or radiation), and unreliable communications continue to limit the usefulness of robots in the field [1].
As robotic technology matures, it becomes increasingly important to ensure its development is informed by the needs and experiences of end-users in the field, especially FRs. Their input is critical in designing systems that can truly support them in the field. In many cases the input of firefighters has been integrated at early stages of the development of these technologies, but in some cases this is not always performed [2, 3].
To help bridge this gap, we conducted a structured survey aimed at gathering feedback from FRs and robot operators across different mission types and backgrounds. Our methodology builds on previous work [4], where we developed a taxonomy of semantics relevant to robotics, and in this study, we presented these concepts to FRs to gather their opinions. We asked them how semantic information, like social media-based Situational Awareness (SA) [5], might support robotic deployments in real-world operations.
@@ II Related work
This work aims at investigating SA, in the context of Search and Rescue (SAR). SA has been defined by Endlsey [6] as three levels - “i) the perception of the elements in the environment within a volume of time and space, ii) the comprehension of their meaning, and iii) the projection of their status in the near future”.
SA is a well-defined concept that encounters significant challenges whenever researchers try to measure the level acquired during operations [7]. The different techniques can be direct or indirect to measure different aspects of SA. This highlights a challenge in objectively determining whether a product guarantees an SA level that is adequate for the task that is to be completed. When talking about dynamic SA in the scope of human-robot teams, [8] systematically discussed the dynamic nature of team SA and the factors that affect SA.
In disaster management, the problem of acquiring SA has been analyzed by several researchers. Some works propose solutions to improve SA in different applications. For example, the work from Allison et al. [9] proposes a framework to control different robots in a hazardous environment with the objective of collecting information and improving the decision-making process. Similarly, in the work proposed by Yang et al. [10], the goal was to develop a robot able to aid in the SAR task by increasing the information collected, and consequently the SA acquired during the mission. In both of these works, the focus is on the engineering side of the solution and not on how the solution improves the SA of the end users. The effect on the SA is not measured, and the opinion of the end users has not been formally collected.
Other works, such as [11] and [12], focus more on the human aspect of the task. The work detailed in [11] focuses on the requirements of FRs for robotic deployment in a SAR scenario. On the other hand, the work from Murphy et al. [12] focuses on the collection of the challenges in rural areas for what concerns disaster management operations. Both of these works propose a series of interviews with experts in the field to gather information from FR experts. In a similar way, with our work, we propose to investigate how FRs acquire SA and if they would be open to using robots with semantic information to improve the decision-making process.
@@ III Methodology
In the questionnaires, we investigate the FRs’ backgrounds from two key dimensions: personal information and experiences with robots. The demographic characteristics include the following features: age, gender, country, education level, roles in the team/field, experience in the field, experience with robots, and others. Next, we propose a validated questionnaire, General Attitudes Towards Robots (GAToRS) [13], as they may affect how FRs perceive domain-specific robotics. The Italian version of the GAToRS has been translated following the validated translation provided by Carradore et al. [14]. In the questionnaire section Situational awareness and use of semantic information, we survey FRs about their usage, experience, expectations, and overall opinions on SA and semantics.
@@ IV-B General attitudes towards robots
The general attitude section is designed based on [13]. The validated questionnaire is the General Attitudes Towards Robots Scale (GAToRS). The scale differentiates i) comfort and enjoyment around robots, ii) unease and anxiety around robots, iii) rational hopes about robots in general (at societal level), and iv) rational worries about robots in general (at societal level) [13]. Hence, the questionnaire is divided into four dimensions:
@@ IV-C Views about SA and use of semantic information
These questions are answered on a Likert scale [15] from one to five, where one refers to a low level (e.g. totally disagree, low frequency etc.), and five refers to a high level (totally agree, high frequency etc.).
@@ V Analysis and discussion
The challenges to obtain SA (Q45) come from different aspects of field deployments, including hardware and device malfunction [16], complex environments, lacking adequate perception, disorganized data, and communication problems with team members.
Based on the feedback of (Q46), we summarize the important semantics in the mission in Fig.˜8. If we fit these semantics into the proposed semantic taxonomy in [4], we find many of them are low-level semantics that can be obtained by robots e.g. location, distance in the blue box. Some of them are high-level semantics e.g. risk in the red box. And many of them are context e.g. medical center, emergency infrastructure in the green box. According to the answers, Junior FRs, who work on the front line and handle the specific operations, focus on the semantics in the blue and red boxes. Senior FRs, who are often the team leader or commander, need to do overall management of the mission. They focus on the semantics and context in the red and green boxes more often.
Results from Q47 and Q48 show that most FR experienced a slight loss of concentration, fatigue, and cognitive overload during the mission (mean 2.8 on a score from one to five, median 3, see Fig.˜3). The sources of fatigue and cognitive overload are tasks, environments, and ethical professional constraints. This problem will not be solved by the application of a specific technology, but can be partially alleviated by the use of semantics [17].

### v2

@@ Abstract
To the best of our knowledge, this study offers novel insights by being one of the first to directly survey FRs on semantic-based SA in a cross-national context. It reveals the types of semantic information most valued in the field-such as object identity, spatial relationships, and risk context-and connects these preferences to the respondents’ roles, experience, and education levels. The findings also expose a critical gap between lab-based robotics capabilities and the realities of field deployment, highlighting the need for more meaningful collaboration between FRs and robotics researchers. These insights contribute to the development of more user-aligned and situationally aware robotic systems for emergency response.
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
At the same time, robotic technologies — such as drones, ground vehicles, and other autonomous systems — are becoming increasingly accessible and capable, due to rapid advances in perception, planning, and control. Commercial applications are expanding, yet these developments have not fully translated into effective tools for FRs operating in real-world extreme conditions. Despite growing interest, a significant gap persists between laboratory robotics and field deployments. Challenges such as rugged terrain, environmental hazards (e.g., heat or radiation), and unreliable communications continue to limit the usefulness of robots in the field [huang2023went].
As robotic technology matures, it becomes increasingly important to ensure its development is informed by the needs and experiences of end-users in the field, especially FRs. Their input is critical in designing systems that can truly support them in the field. In many cases the input of firefighters has been integrated at early stages of the development of these technologies, but in some cases this is not always performed [10558013, mcconville2024adoption].
To help bridge this gap, we conducted a structured survey aimed at gathering feedback from FRs and robot operators across different mission types and backgrounds. Our methodology builds on previous work [ruan2022taxonomy], where we developed a taxonomy of semantics relevant to robotics, and in this study, we presented these concepts to FRs to gather their opinions. We asked them how semantic information, like social media-based Situational Awareness (SA) [snyder2019situational], might support robotic deployments in real-world operations.
@@ II Related work
This work aims at investigating SA, in the context of Search and Rescue (SAR). SA has been defined by Endlsey [endlsey1995] as three levels - “i) the perception of the elements in the environment within a volume of time and space, ii) the comprehension of their meaning, and iii) the projection of their status in the near future”.
SA is a well-defined concept that encounters significant challenges whenever researchers try to measure the level acquired during operations [Cooper2013-bq]. The different techniques can be direct or indirect to measure different aspects of SA. This highlights a challenge in objectively determining whether a product guarantees an SA level that is adequate for the task that is to be completed. When talking about dynamic SA in the scope of human-robot teams, [senaratne2025framework] systematically discussed the dynamic nature of team SA and the factors that affect SA.
In disaster management, the problem of acquiring SA has been analyzed by several researchers. Some works propose solutions to improve SA in different applications. For example, the work from Allison et al. [allison2024] proposes a framework to control different robots in a hazardous environment with the objective of collecting information and improving the decision-making process. Similarly, in the work proposed by Yang et al. [yang2014], the goal was to develop a robot able to aid in the SAR task by increasing the information collected, and consequently the SA acquired during the mission. In both of these works, the focus is on the engineering side of the solution and not on how the solution improves the SA of the end users. The effect on the SA is not measured, and the opinion of the end users has not been formally collected.
Other works, such as [betta2024perceptions] and [murphy2025rural], focus more on the human aspect of the task. The work detailed in [betta2024perceptions] focuses on the requirements of FRs for robotic deployment in a SAR scenario. On the other hand, the work from Murphy et al. [murphy2025rural] focuses on the collection of the challenges in rural areas for what concerns disaster management operations. Both of these works propose a series of interviews with experts in the field to gather information from FR experts. In a similar way, with our work, we propose to investigate how FRs acquire SA and if they would be open to using robots with semantic information to improve the decision-making process.
@@ III Methodology
In the questionnaires, we investigate the FRs’ backgrounds from two key dimensions: personal information and experiences with robots. The demographic characteristics include the following features: age, gender, country, education level, roles in the team/field, experience in the field, experience with robots, and others. Next, we propose a validated questionnaire, General Attitudes Towards Robots (GAToRS) [koverola2022general], as they may affect how FRs perceive domain-specific robotics. The Italian version of the GAToRS has been translated following the validated translation provided by Carradore et al. [Carradore2024]. In the questionnaire section Situational awareness and use of semantic information, we survey FRs about their usage, experience, expectations, and overall opinions on SA and semantics.
@@ IV-B General attitudes towards robots
The general attitude section is designed based on [koverola2022general]. The validated questionnaire is the General Attitudes Towards Robots Scale (GAToRS). The scale differentiates i) comfort and enjoyment around robots, ii) unease and anxiety around robots, iii) rational hopes about robots in general (at societal level), and iv) rational worries about robots in general (at societal level) [koverola2022general]. Hence, the questionnaire is divided into four dimensions:
@@ IV-C Views about SA and use of semantic information
These questions are answered on a Likert scale [joshi2015likert] from one to five, where one refers to a low level (e.g. totally disagree, low frequency etc.), and five refers to a high level (totally agree, high frequency etc.).
@@ V Analysis and discussion
The challenges to obtain SA (Q45) come from different aspects of field deployments, including hardware and device malfunction [ramachandran2021resilient], complex environments, lacking adequate perception, disorganized data, and communication problems with team members.
Based on the feedback of (Q46), we summarize the important semantics in the mission in Fig.˜8. If we fit these semantics into the proposed semantic taxonomy in [ruan2022taxonomy], we find many of them are low-level semantics that can be obtained by robots e.g. location, distance in the blue box. Some of them are high-level semantics e.g. risk in the red box. And many of them are context e.g. medical center, emergency infrastructure in the green box. According to the answers, Junior FRs, who work on the front line and handle the specific operations, focus on the semantics in the blue and red boxes. Senior FRs, who are often the team leader or commander, need to do overall management of the mission. They focus on the semantics and context in the red and green boxes more often.
Results from Q47 and Q48 show that most FR experienced a slight loss of concentration, fatigue, and cognitive overload during the mission (mean 2.8 on a score from one to five, median 3, see Fig.˜3). The sources of fatigue and cognitive overload are tasks, environments, and ethical professional constraints. This problem will not be solved by the application of a specific technology, but can be partially alleviated by the use of semantics [ruan2025exploratory].